# Tuning Language Models for Robust Prediction of Diverse User Behaviors

### v1

@@ 4.3.1 Performance robustness
In Figure 3(a), We evaluate BehaviorLM on anchor and tail behaviors (Behavior dataset) across model sizes and tuning strategies. Scaling from 1.5B to 70B significantly improves accuracy—for anchors (0.415 → 0.451) and tails (0.354 → 0.385)—demonstrating robustness. However, without our progressive tuning (A-Tuning + B-Tuning), larger models yield only marginal gains (blue curve in Figure 3(a)), highlighting the effectiveness of our approach in leveraging LLM behavioral knowledge.
@@
Additionally, we conduct efficiency experiments demonstrating that the behavioral knowledge in LLMs offers over 100× greater sample efficiency than traditional transformers. Experimental details are in the E.
4.4 Ablation Study
@@
In Table 5, we evaluate the contribution of each design component to overall performance through ablation studies on the Behavior dataset. Specifically, we examine the performance drop when: (1) removing the auxiliary task in A-tuning, (2) replacing difficulty-based data selection with uniform selection, (3) performing non-contrastive data selection by setting the distinguishability score , and (4)replacing A-tuning with fine-tuning on all behaviors, as done in previous works Bao et al. (2023); Liao et al. (2024). The ablation experiment results of the APP dataset are detailed in the Appendix F.
Key findings are as follows: All design components enhance prediction performance. For head behaviors, B-tuning’s data selection is most impactful by improving behavior differentiation. For medium and tail behaviors, A-tuning is essential. Without it, accuracy drops sharply, as direct fine-tuning dilutes performance on tail behaviors.
Dataset | BehaviorLM Variant | Overall | Head | Medium | Tail |
Behavior | A-Tuning w/o Aux. Task | -4.71% | -1.07% | -5.99% | -4.58% |
B-Tuning w/o DDS | -6.23% | -4.34% | -7.66% | -5.62% | |
B-Tuning w/o CDS | -5.50% | -3.24% | -6.83% | -5.02% | |
w/o A-Tuning | -10.00% | -2.77% | -11.74% | -10.61% |
4.5 Investigating the Necessity of Anchor Behavior Tuning
@@
5 related works
5.1 LLM-Enhanced Behavior Prediction Models
@@
User behavior prediction based on the most recent events is similar to sequential recommendation. The key difference is that behavior prediction focuses on recurring daily actions, while item recommendation emphasizes novel content. Due to limited research in behavior prediction, we draw on related work in sequential recommendation. Recent studies in recommendation explore knowledge alignment between language and recommendation domains. For example, A-LLMRec used multi-task training for embedding alignment Kim et al. (2024), PLM-Rec applied mutual information maximizationGeng et al. (2022) , PITuning introduced PITuning for cross-modal pattern extractionGong et al. (2024), and LLM-ESR proposed a cross-attention mechanism for sequence alignmentLiu et al. (2024). While these methods enhance traditional models by aligning language and recommendation knowledge, they underutilize LLMs’ zero-shot and few-shot generalization. Our approach addresses this gap with progressive fine-tuning, preserving general behavioral knowledge while improving prediction of infrequent long-tail behaviors without sacrificing performance on frequent ones.
5.2 LLM-Based Behavior Prediction Models
Traditional behavior prediction models often rely on embedding-based neural networks, which lack interpretability. LLMs, with their rich world knowledge and reasoning capabilities, present a more interpretable alternative for user behavior prediction. Tallrec first showcased LLMs’ strong few-shot performance in recommendation tasks Bao et al. (2023), while Llara improved fine-tuning through curriculum learning, progressing from simpler to more complex tasks Liao et al. (2024). Similarly, Recexplainer aligned LLMs with traditional recommendation models for user-specific, customizable explanations Lei et al. (2024) . Other notable frameworks include the Chain-of-Planned Behavior Shao et al. (2024), which captured spatial-temporal dynamics, and the Transition paradigm TransRec Lin et al. (2024), combining multiple identifiers to enhance LLM-based recommendations. However, these approaches mainly convert behavior sequences into text for LLM training, often overlooking critical distinctions between anchor and tail behaviors, which hampers their zero-shot generalization. Our method, BehaviorLM, directly addresses the long-tailed behavior distribution by employing progressive fine-tuning. This two-stage process significantly improves the prediction of tail behaviors without compromising anchor behavior performance.
6 Conclusion
@@
In this paper, we leverage the rich behavioral knowledge in LLMs to tackle user behavior prediction, with a focus on long-tail behavior prediction. We propose a progressive tuning approach, where the LLM first learns frequent anchor behaviors before generalizing to rarer tail behaviors. Experiments on two real-world datasets show that BehaviorLM outperforms state-of-the-art methods, achieving up to 27%/20% improvement in long-tail behavior prediction, addressing a traditionally challenging aspect of behavior modeling. Analysis highlights that addressing the long-tailed behavior distribution is essential for effectively utilizing LLMs’ behavioral knowledge in fine-tuning.
@@ Appendix A Details of Used Datasets
Behavior dataset: This large-scale dataset is derived from mobile phone usage logs. When users interact with their mobile phones, various types of logs are generated, desensitized, and reported with user consent. We select 37 daily behaviors that are reliably extracted from raw logs and also cover broad life scenarios, including activities related to learning, work, entertainment, leisure, etc. The dataset spans from March 1, 2024, to April 29, 2024, and consists of over 50 million behavior events from 24,133 anonymous users. We preprocess the dataset and construct samples in the format of “(last-20 events, next event)”. Since our target is fine-tuning the LLM instead of training from scratch, we only randomly select a subset (200,000) of data for experiments.
App usage dataset [16]: This dataset is an open-source resource that captures 1753 user interactions across various apps within one week. Given the large number of apps and the overlap in functionality among many of them, we processed the dataset further by merging apps with similar purposes. For example, Douyin and Kuaishou were grouped into a "watching videos" category. After similar preprocessing as the Behavior dataset, this dataset consists of 71 behavior events belonging to 24 categories.
Behavior | Proportion |
Watching video clip | 0.435981 |
Shopping online | 0.112217 |
Editing video | 0.05937 |
Reading books | 0.055293 |
Gaming | 0.044202 |
Watching video | 0.041266 |
Reading news | 0.036046 |
Taking photos | 0.035068 |
Taking subway | 0.031969 |
Payment | 0.029848 |
Listening Music | 0.02365 |
Checking Weather | 0.016963 |
Studying for course | 0.01419 |
Ordering takeout | 0.012722 |
Listening podcast | 0.012396 |
Conducting job recruitment | 0.010928 |
Taking notes | 0.003915 |
Checking calendar | 0.003591 |
Hailing taxi | 0.003262 |
Picking up the express delivery | 0.00261 |
Watching Match | 0.00213 |
Exercising | 0.002003 |
Singing | 0.001957 |
Watching Live-streaming | 0.001794 |
Cycling | 0.001794 |
Meeting online | 0.000816 |
Studying for driving test | 0.000652 |
Translating | 0.000652 |
Rentaling house | 0.000652 |
Watching animation | 0.000489 |
Booking ticket | 0.000489 |
Commuting check-in | 0.000323 |
Cooking | 0.000323 |
Drinking coffee | 0.000283 |
Rentaling car | 0.000163 |
The characteristics of both two datasets are presented in Table 4.
Dataset | # User | # Behavior Type | # Sample |
Behavior Dataset | 24,133 | 37 | 200,000 |
App Usage Dataset | 1,753 | 24 | 50,000 |
@@ Appendix C Baselines
SASRec [12]. uses self-attention mechanisms to model user behavior sequences. It captures both short-term and long-term dependencies in sequential data, allowing it to focus on the most relevant items in the user’s interaction history for recommendation.
BERT4Rec [30]. models user behavior sequences using deep bidirectional self-attention. By jointly considering the context before and after an item, it predicts the randomly masked items within the sequence, achieving excellent predictive performance.
LLaRa [17] introduces a hybrid prompting method that integrates both world knowledge and behavioral patterns into item representations. It conducts curriculum prompt tuning to achieve modality alignment.
A-LLMRec [13] bridges the knowledge between the language and recommendation domains by training an alignment network with a variety of tasks, targeting both warm and cold-start scenarios.
PITuning [9] loads pre-trained Large Language Model (LLM) parameters to acquire textual knowledge and then designs an adaptive unlearning strategy to address the long-tail preference issue, achieving excellent performance in user behavior prediction.
LLM-ESR [19] leverages the semantic information from LLMs, proposes a dual-view modeling framework enhanced through embedding techniques to capture the nuances of long-tail items better, demonstrating strong performance across multiple datasets..
TALLRec [3] is one of the earlier methods to integrate Large Language Models (LLMs) with the recommendation domain. It employs a two-stage tuning process—Alpaca Tuning and Rec-Tuning—to finetune LLMs for recommendations, enabling effective and efficient adaptation of LLMs with only a small number of tuning samples.
AlphaFuse [11] is a simple yet effective language-guided learning strategy that addresses long-tail intent modeling by learning ID embeddings within the null space of language embeddings.
Appendix D Implementation Details
The hardware used in this experiment consists of 8 NVIDIA A100 40G GPUs. We selected LLama3.1-8B [7] as the backbone for our experiments. To ensure flexibility in model testing, we designed three distinct instruction formats, which are randomly sampled during both training and testing. Our experiments utilized the AdamW optimizer with a cosine annealing learning rate schedule, setting the warm-up proportion to 0.1. The maximum learning rate for cosine annealing was set to 1e-4, while both the minimum and initial warm-up learning rates were set to 1e-6. We conducted LoRA fine-tuning and parallel training acceleration using the open-source LLM instruction fine-tuning library, llama-factory [38]. All experiments were performed with a maximum of 8 training epochs and a batch size of 8, selecting the best-performing model on the validation set for testing. Detailed formats of the three instruction types are as followed:

### v2

@@ 4.3. Investigating the Effect of Behavioral Knowledge
In Figure 3(a), we compare BehaviorLM’s prediction performance on anchor and tail behaviors (Behavior dataset) using different model sizes and tuning strategies. Increasing the model size from 1.5B to 70B significantly improves behavior prediction accuracy. Specifically, the relative improvement is similar for anchor behaviors and tail behaviors, demonstrating prediction robustness. Furthermore, when fine-tuning the LLM directly on all behaviors—without using our proposed progressive tuning approach (A-Tuning + B-Tuning)—the improvement from increasing model size is relatively marginal (blue curve in the figure). This suggests that our method effectively leverages the behavioral knowledge embedded in the LLM.
@@
4.3.2. Sample efficiency.
One significant advantage of leveraging the LLM’s general behavioral knowledge is that it reduces the need for fine-tuning on large-scale user behavior data, demonstrating strong sample efficiency. To validate this, we compare BehaviorLM-1.5B and BehaviorLM-8B with another transformer-based model trained from scratch, using the objective function from SASRec (Kang and McAuley, 2018) and the model architecture from GPT2 (Radford et al., 2019) (the small version with 12 layers and 768 latent dimensions). Since the original Behavior dataset contains over 50 million events, we vary the sample size to observe its impact on prediction performance. As shown in Figure 4, BehaviorLM demonstrates a significant improvement in sample efficiency. The transformer-based model trained from scratch only outperforms BehaviorLM when trained on nearly all 50 million samples, while BehaviorLM is fine-tuned on just 200,000 samples—over two orders of magnitude fewer. This highlights the remarkable sample efficiency advantage provided by the LLM’s preexisting behavioral knowledge. Using 8xA100 (40G) GPUs, training BehaviorLM-8B on Behavior Dataset with 200k samples takes about 6 hours, while on APP Usage dataset (50000 samples) it takes no more than 2 hours. In contrast, training a SASRec model on a 50M sample dataset requires approximately 72 hours. Overall, although BehaviorLM has a lower training speed for each sample, it has a much higher sample efficiency than SASRec and thus takes much less time to achieve a good prediction accuracy.
4.4. Ablation Study
@@
In Table 4, we evaluate the contribution of each design component to the overall performance through comprehensive ablation studies on both the App Usage and Behavior datasets. Specifically, we examine the performance drop when: (1) w/o Aux. Task: Removing the auxiliary conversation task during A-tuning, (2) w/o DDS: Replacing the difficulty-based data selection with uniform random selection during B-Tuning, (3) w/o KMeans: Removing the semantic clustering step (Difficulty-Weighted K-Means++), (4) w/o DPO: Replacing the Direct Preference Optimization in B-Tuning with standard Supervised Fine-Tuning (SFT), and (5) w/o A-Tuning: Skipping the first stage and fine-tuning the LLM directly on all behavior types.
| (a) App Usage Dataset | ||||
| BehaviorLM Variant | Overall | Head | Medium | Tail |
| A-Tuning w/o Aux. Task | -3.91% | -2.45% | -2.19% | -6.00% |
| B-Tuning w/o DDS | -7.90% | -5.74% | -6.30% | -10.11% |
| B-Tuning w/o Keans | -2.76% | -1.60% | -2.95% | -3.52% |
| B-Tuning w/o DPO | -2.55% | -2.31% | -1.32% | -2.61% |
| w/o A-Tuning | -8.03% | -0.93% | -5.87% | -14.26% |
| (b) Behavior Dataset | ||||
| BehaviorLM Variant | Overall | Head | Medium | Tail |
| A-Tuning w/o Aux. Task | -4.57% | -1.10% | -6.11% | -4.40% |
| B-Tuning w/o DDS | -6.29% | -4.56% | -7.51% | -5.56% |
| B-Tuning w/o Keans | -3.08% | -0.90% | -4.43% | -4.65% |
| B-Tuning w/o DPO | -3.58% | -2.84% | -5.61% | -1.74% |
| w/o A-Tuning | -9.80% | -2.83% | -11.39% | -10.93% |
The key conclusions from Table 4 are as follows: First, all the design components contribute to the final prediction performance. For head-category behaviors, the difficulty-based data selection strategy in B-tuning has the most significant impact, as it improves the model’s ability to differentiate between different behaviors, avoiding the erroneous and significant suppression of head-category behavior predictions. For medium and tail categories, the largest contribution comes from the A-tuning stage, without which prediction accuracy drops significantly. This is expected, as fine-tuning the LLM directly on all behaviors causes a loss in predictive capability for tail behaviors, which cannot be recovered through few-shot tuning in the B-tuning stage. Furthermore, w/o KMeans shows a moderate decline, validating that while difficulty is crucial, ensuring semantic diversity through clustering prevents the model from overfitting to specific types of failure cases. Replacing DPO with SFT (w/o DPO) results in a consistent performance drop, suggesting that the contrastive nature of DPO better leverages the hard negatives generated during the difficulty scoring phase. Finally, removing the auxiliary task (w/o Aux. Task) causes a moderate decline, proving that maintaining general capabilities via multi-task learning acts as a necessary regularizer to support robust behavior prediction.
4.5. Investigating the Necessity of Anchor Behavior Tuning
@@
Additionally, we conducted hyperparameter experiments on the Behavior dataset to assess the impact of varying auxiliary task data proportions. As evident from Table 5, an appropriate amount of auxiliary task data can significantly enhance the model’s performance. Insufficient auxiliary task data fails to deliver notable improvements, whereas an excessive amount can disrupt the model’s predictive capabilities.
| Auxiliary Task Data Ratio | overall | head | medium | tail |
| 0% | 0.404 | 0.556 | 0.418 | 0.376 |
| 2.5% | 0.423 | 0.543 | 0.432 | 0.386 |
| 5% | 0.439 | 0.556 | 0.452 | 0.398 |
| 10% | 0.416 | 0.542 | 0.423 | 0.382 |
| 20% | 0.410 | 0.522 | 0.414 | 0.348 |
5. Extended Hyperparameter and Mechanism Analysis
In this section, we provide a deeper analysis regarding the hyperparameters used for data partition and investigate the underlying mechanism of the auxiliary task utilized in the A-Tuning stage.
5.1. Sensitivity to Anchor/Tail Threshold
In our main experiments, we defined “Anchor Behaviors” as those occurring with a frequency greater than 1%. To evaluate the robustness of BehaviorLM with respect to this threshold selection, we conducted additional experiments on the Behavior Dataset. We varied the threshold for distinguishing anchor and tail behaviors to 0.5%, 1% (the setting used in the main paper), and 2%.
The results are presented in Table 6. As observed, BehaviorLM achieves consistent performance across different threshold settings. For instance, the Overall Accuracy remains stable around 0.42, and the Tail Accuracy shows only minor fluctuations. This demonstrates that our method is robust to the specific choice of the long-tail threshold and does not rely on a specific data partition to achieve superior performance.
| Threshold | ||||||
| 0.5% | 0.6301 | 0.6429 | 0.4449 | 0.5394 | 0.4382 | 0.4202 |
| 1.0% (Default) | 0.6317 | 0.6402 | 0.4389 | 0.5558 | 0.4506 | 0.3978 |
| 2.0% | 0.6250 | 0.6368 | 0.4322 | 0.5445 | 0.4304 | 0.3973 |
5.2. Investigation into the Mechanism of Auxiliary Tasks
To improve the model’s generalization capability during the A-Tuning stage, we incorporated a general conversational dataset (ShareGPT) as an auxiliary task. A key question arises regarding the mechanism of this improvement: does it stem from specific knowledge transfer or a general regularization effect? We hypothesize that the auxiliary dialogue task functions primarily through a regularization effect. By maintaining the LLM’s general text generation capabilities, it prevents overfitting to the anchor behavior prediction task and mitigates the catastrophic forgetting of common sense.
To validate this hypothesis, we performed experiments on the Behavior Dataset replacing the original ShareGPT data with two alternative domains:
MATH (Hendrycks et al., 2021): A dataset focused on mathematical reasoning problems.
HealthQA (Hosseini et al., 2024): A medical dialogue dataset that is semantically unrelated to the user behavior domain.
The results, summarized in Table 7, show that model performance does not differ significantly across the three diverse auxiliary tasks. Whether using general chat, mathematical reasoning, or medical dialogue, the improvement in tail behavior prediction remains comparable. This reinforces our conclusion that the benefit of the auxiliary task stems from general regularization—helping the model maintain its intrinsic reasoning and generation abilities—rather than domain-specific semantic knowledge transfer.
| Auxiliary Task Type | ||||||
| ShareGPT (Default) | 0.6317 | 0.6402 | 0.4389 | 0.5558 | 0.4506 | 0.3978 |
| MATH | 0.6293 | 0.6344 | 0.4451 | 0.5471 | 0.4328 | 0.4224 |
| HealthQA | 0.6330 | 0.6355 | 0.4517 | 0.5481 | 0.4448 | 0.4267 |
5.3. Generalization and Robustness
To assess the generalization capability of BehaviorLM on long-tailed tasks beyond smartphone usage, we extended our evaluation to the Foursquare-NYC dataset (Yang et al., 2014), a representative benchmark in the mobility prediction domain. Furthermore, we examined the robustness of our method with respect to hyperparameter settings. It is worth noting that we employed the identical configuration for the Foursquare-NYC dataset as used in the main experiments on App Usage and Behavior datasets, without performing separate hyperparameter optimization. Specifically, we fixed the anchor threshold at 1%, the auxiliary task ratio at 5%, and the number of samples per behavior type for B-Tuning at 20.
As presented in Table 8, BehaviorLM consistently outperforms all baselines across all metrics. Notably, in the tail segment, our method achieves a Tail Accuracy of 0.2157, significantly surpassing the best-performing baseline (GPT4o with 0.1405), confirming that our progressive tuning strategy effectively generalizes to other long-tailed domains. And the consistent performance achieved under these fixed settings across three diverse datasets demonstrates that BehaviorLM is highly robust and does not rely on extensive, dataset-specific hyperparameter tuning.
| Model | Overall | Head | Medium | Tail | ||
| SASRec | 0.2781 | 0.2841 | 0.1307 | 0.2202 | 0.0980 | 0.0065 |
| Bert4Rec | 0.2080 | 0.2445 | 0.0556 | 0.1205 | 0.0000 | 0.0000 |
| PITuning | 0.3243 | 0.3167 | 0.1220 | 0.2132 | 0.1225 | 0.0000 |
| LLM-ESR | 0.2834 | 0.2896 | 0.1430 | 0.2238 | 0.1085 | 0.0152 |
| AlphaFuse | 0.3313 | 0.3217 | 0.1463 | 0.2246 | 0.1493 | 0.0240 |
| LLama-NT | 0.2822 | 0.2826 | 0.1514 | 0.1863 | 0.1372 | 0.1144 |
| GPT4o | 0.2989 | 0.3010 | 0.1912 | 0.2273 | 0.2056 | 0.1405 |
| A-LLMRec | 0.2819 | 0.2769 | 0.0880 | 0.1520 | 0.0932 | 0.0882 |
| LLaRA | 0.3248 | 0.3245 | 0.1786 | 0.2206 | 0.2059 | 0.1046 |
| CoLLM | 0.3287 | 0.3196 | 0.1824 | 0.2157 | 0.2069 | 0.1075 |
| TALLRec | 0.3246 | 0.3220 | 0.1786 | 0.2181 | 0.2157 | 0.1013 |
| BehaviorLM | 0.3508 | 0.3471 | 0.2386 | 0.2525 | 0.2451 | 0.2157 |
6. Related Works
6.1. LLM-Enhanced Behavior Prediction Models
@@
User behavior prediction based on the most recent events is similar to sequential recommendation. The key difference is that behavior prediction focuses on recurring daily actions, while item recommendation emphasizes novel content. Due to limited research in behavior prediction, we draw on related work in sequential recommendation. Recent studies in recommendation explore knowledge alignment between language and recommendation domains. For example, SAID(Hu et al., 2024) proposed improving sequential recommendation via LLM-based semantic embedding learning, PLM-Rec applied mutual information maximization(Geng et al., 2022) , PITuning introduced PITuning for cross-modal pattern extraction(Gong et al., 2024), and LLM-ESR proposed a cross-attention mechanism for sequence alignment(Liu et al., 2024). In addition, to address the issue of sparse user data, ProEx(Zhang et al., 2025c) utilizes LLMs for profile extrapolation to generate enriched user representations, augmenting the input for downstream recommendation models. LLMCDSR(Xin et al., 2025) leverages LLMs to transfer knowledge and enhance user interest modeling in sparse domains, effectively mitigating the cold-start problem for traditional predictors. While these methods enhance traditional models by aligning language and recommendation knowledge, they underutilize LLMs’ zero-shot and few-shot generalization. Our approach addresses this gap with progressive fine-tuning, preserving general behavioral knowledge while improving prediction of infrequent long-tail behaviors without sacrificing performance on frequent ones.
6.2. LLM-Based Behavior Prediction Models
Current behavior prediction models largely rely on embedding neural networks, which often function as black-box models with limited interpretability. LLMs, with their vast world knowledge and powerful reasoning abilities, offer an alternative approach that enhances interpretability in user behavior prediction (Wu et al., 2024b). The introduction of LLMs into behavior prediction was pioneered by Bao et al. (2023), who demonstrated the impressive few-shot performance of LLMs in the recommendation domain. To formalize this paradigm, Zhang et al. (2025a) proposed ”Recommendation as Instruction Following,” establishing a unified framework where the LLM directly outputs items based on natural language instructions. Building on these foundations, several strategies have been proposed to optimize LLM performance. Liao et al. (2024) employed a curriculum learning strategy to progressively fine-tune LLMs from easier tasks to more complex ones, while Kim et al. (2024) utilized multi-task training for embedding alignment. Furthermore, Mao et al. (2025) applied reinforcement learning to dynamically optimize prompts, enhancing the LLM’s ability to capture personalized user intent. Regarding complex data distributions, Wu et al. (2024a) introduced CoRAL, a retrieval-augmented LLM framework that explicitly improves long-tail recommendation by retrieving collaborative evidence. Similarly, Xia et al. (2025) proposed a hierarchical tree search mechanism based on LLMs to model lifelong user behaviors. Other approaches focus on specific modeling aspects. For instance, the Chain-of-Planned Behavior framework by Shao et al. (2024) captures the spatial-temporal dynamics of user activities. Lin et al. (2024) introduced a Transition paradigm combining multiple identifiers to enhance recommendations, and Lei et al. (2024) proposed alignment techniques to train LLMs to mimic traditional recommender models for customizable explanations.
However, these methods primarily focus on converting behavior sequences into textual representations for LLM training or enhancing external information retrieval, often neglecting the critical differences between anchor and tail behaviors. This limits their zero-shot generalization ability. Our proposed approach, BehaviorLM, differentiates them by addressing the long-tailed distribution of behaviors. First, we fine-tune on anchor behaviors while preserving general behavioral knowledge, and second, we fine-tune on a balanced subset of behaviors based on sample difficulty, significantly enhancing the prediction of tail behaviors without sacrificing anchor behavior performance.
7. Conclusion
@@
In this paper, we leverage the rich behavioral knowledge in LLMs to tackle user behavior prediction, with a focus on long-tail behavior prediction. We propose a progressive tuning approach, where the LLM first learns frequent anchor behaviors before generalizing to rarer tail behaviors. Experiments on two real-world datasets show that BehaviorLM outperforms state-of-the-art methods, achieving up to 30.8%/22.5% improvement in long-tail behavior prediction, addressing a traditionally challenging aspect of behavior modeling. Analysis highlights that addressing the long-tailed behavior distribution is essential for effectively utilizing LLMs’ behavioral knowledge in fine-tuning.
@@ Appendix A Details of Used Datasets
Detailed statistics of the Behavior Dataset are shown in the figure 6.
@@ Appendix C Baselines
SASRec (Kang and McAuley, 2018). uses self-attention mechanisms to model user behavior sequences. It captures both short-term and long-term dependencies in sequential data, allowing it to focus on the most relevant items in the user’s interaction history for recommendation.
BERT4Rec (Sun et al., 2019). models user behavior sequences using deep bidirectional self-attention. By jointly considering the context before and after an item, it predicts the randomly masked items within the sequence, achieving excellent predictive performance.
CoLLM (Zhang et al., 2025b) captures collaboration information using external traditional models and maps it into the LLM’s input embedding space as collaboration embeddings. This external integration allows effective modeling of collaboration without modifying the LLM, enabling flexible use of various collaboration modeling techniques.
LLaRa (Liao et al., 2024) introduces a hybrid prompting method that integrates both world knowledge and behavioral patterns into item representations. It conducts curriculum prompt tuning to achieve modality alignment.
A-LLMRec (Kim et al., 2024) bridges the knowledge between the language and recommendation domains by training an alignment network with a variety of tasks, targeting both warm and cold-start scenarios.
PITuning (Gong et al., 2024) loads pre-trained Large Language Model (LLM) parameters to acquire textual knowledge and then designs an adaptive unlearning strategy to address the long-tail preference issue, achieving excellent performance in user behavior prediction.
LLM-ESR (Liu et al., 2024) leverages the semantic information from LLMs, proposes a dual-view modeling framework enhanced through embedding techniques to capture the nuances of long-tail items better, demonstrating strong performance across multiple datasets..
TALLRec (Bao et al., 2023) is one of the earlier methods to integrate Large Language Models (LLMs) with the recommendation domain. It employs a two-stage tuning process—Alpaca Tuning and Rec-Tuning—to finetune LLMs for recommendations, enabling effective and efficient adaptation of LLMs with only a small number of tuning samples.
AlphaFuse (Hu et al., 2025) is a simple yet effective language-guided learning strategy that addresses long-tail intent modeling by learning ID embeddings within the null space of language embeddings.
Appendix D Details of Used Instructions